# Towards Boosting LLMs-driven Relevance Modeling with Progressive Retrieved Behavior-augmented Prompting

### v1

@@
Abstract.
Relevance modeling is a critical component for enhancing user experience in search engines, with the primary objective of identifying items that align with users’ queries. Traditional models only rely on the semantic congruence between queries and items to ascertain relevance. However, this approach represents merely one aspect of the relevance judgement, and is insufficient in isolation. Even powerful Large Language Models (LLMs) still cannot accurately judge the relevance of a query and an item from a semantic perspective. To augment LLMs-driven relevance modeling, this study proposes leveraging user interactions recorded in search logs to yield insights into users’ implicit search intentions. The challenge lies in the effective prompting of LLMs to capture dynamic search intentions, which poses several obstacles in real-world relevance scenarios, i.e., the absence of domain-specific knowledge, the inadequacy of an isolated prompt, and the prohibitive costs associated with deploying LLMs. In response, we propose , a novel gressive etrieved ehavior-augmented rompting framework for integrating search scenario-oriented knowledge with LLMs effectively. Specifically, we perform the user-driven behavior neighbors retrieval from the daily search logs to obtain domain-specific knowledge in time, retrieving candidates that users consider to meet their expectations. Then, we guide LLMs for relevance modeling by employing advanced prompting techniques that progressively improve the outputs of the LLMs, followed by a progressive aggregation with comprehensive consideration of diverse aspects. For online serving, we have developed an industrial application framework tailored for the deployment of LLMs in relevance modeling. Experiments on real-world industry data and online A/B testing demonstrate our proposal achieves promising performance.
1. Introduction
@@
To address the above problems, we propose a novel gressive etrieved ehavior-augmented rompting framework for integrating search scenario-oriented knowledge with LLMs (dubbed as ). To acquire domain-specific knowledge in time, we perform a user-driven behavior neighbor retrieval from the daily updated search logs, retrieving candidates that users consider to meet their expectations currently. Then we anticipate employing advanced prompting techniques that progressively improve the outputs of the LLMs, followed by a progressive aggregation with comprehensive consideration of diverse aspects to form a holistic relevance model. As for the online serving of LLMs, we design an industrial implementation framework enabling LLMs to fully handle search relevance scenarios with the affordable cost.
@@
We propose with two novel plug-in modules. Firstly, a user-driven behavior neighbors retrieval is developed to acquire domain-specific knowledge in time. Secondly, the proposal of progressive prompting and aggregation can strengthen the judgement of relevance for LLMs.
We explore an industrial implementation enabling LLMs to fully handle search relevance scenarios in Alipay search with the affordable cost and latency.
We demonstrate framework achieves superior performance through experiments on real-world industry data and online A/B testing. It has been deployed online and outperforms prior approaches (Chen et al., 2023) on core metrics.
2. Related Work
2.1. Relevance Modeling
The section concretely discusses the most recent advances in relevance modeling studies. Relevance modeling in search can be viewed as a text matching problem as the sub-domain of information retrieval (IR). The majority of work focuses on distinctions at the semantic level, while a minority of methods judge the relevance from a behavioral perspective.
2.1.1. Semantics-driven Methods
@@
In recent times, pre-trained models like BERT (Devlin et al., 2018) have made remarkable strides in Natural Language Understanding (NLU). As a result, representation-based (Yao et al., 2022; Lu et al., 2020; Reimers and Gurevych, 2019; Jin et al., 2023) and interaction-based architectures (Wang et al., 2019) are proposed to leverage the capabilities of these models to encode semantic correlations between queries and items. Most recently, Large Languages Models (LLMs) like GPT (Radford et al., 2019), LLaMA (Touvron et al., 2023), BLOOM (Workshop et al., 2022) and GLM (Du et al., 2022) trained on massive corpora of texts have shown their superior ability in language understanding, generation, interaction, and reasoning tasks. (Sun et al., 2023) investigates the potential of utilizing LLMs for searching and demonstrates that appropriately instructs ChatGPT and GPT-4 can produce competitive and even superior results to supervised methods widely used information retrieval benchmarks. (Chen et al., 2023) tries to deal with long-tail query-item matching through LLMs efficiently and effectively. In these research work, they tend to exploit world knowledge stored in parameters of LLMs to judge the relevance between the query and item. However, general LLMs can not adapt to the industrial scenario due to the lack of domain-specific knowledge and insensitivity to the relevance judgement. In this work, we try to explore the LLMs-driven relevance modeling comprehensively in Alipay search engine to meet the above requirements.
2.1.2. Behavior-driven Methods
@@ 3. Problem Formulation
4. Methodology
The architecture of the proposed framework is depicted in Figure 2. It contains two novel plug-in modules: (1) user-driven behavior neighbor retrieval for obtaining domain-specific knowledge in time to meet users’ expectation; (2) progressive prompting and aggregation for improving the sensitivity of LLMs to relevance judgement and stable prediction results. Through the two novel plug-in modules, relevance degree generation can be performed. In what follows, we elaborate on the two modules.
4.1. User-driven Behavior Neighbor Retrieval
Generally, LLMs possess an exhaustive world knowledge with the benefit of massive corpora of texts pre-training. However, LLMs still struggle to understand user search intentions for short and ambiguous queries and items from Alipay search due to the lack of specialized and rapidly evolving domain knowledge. There are a few approaches could deal with the above problem, e.g., continual pre-training and retrieval-augmented generation. Considering efficiency and cost issues, the promising approach is retrieval-augmented language modeling (Gao et al., 2023), grounding the LLMs during generation by conditioning on relevant candidates retrieved from an external knowledge source. Drawing inspiration from this, we devise a user-driven behavior neighbor retrieval module. This module could retrieve the daily search logs of users to obtain daily changing behavior neighbors that users consider relevant. The pipeline of user-driven behavior neighbor retrieval is depicted as Figure 1.
@@ 4.1. User-driven Behavior Neighbor Retrieval
Due to the limited number of items displayed for a query in Alipay search scenario, almost all results can be seen by users. With this in mind, we could analyze that a higher click-through rate reflects that users believe the corresponding query-item pairs can better meet their search intents and needs when the exposure PV (i.e., page view) reaches a certain quantity. Thus, we filter out the query-item pairs with less than 100 exposure PV and formulate the remaining search logs as high-confidence logs. Then, we utilize the high-confidence logs from the past month to calculate the click-through rate for the exposed query-item pairs. To mitigate the effect of noises, query-item pairs are selected above a click-through threshold (e.g., 0.2) and the neighbors are arranged in descending order based on click-through rate from the query and item perspective, respectively. And we only choose top-K (e.g., 20) neighbors for corresponding query and item. Through this approach, it is possible to select dual behavior neighbors with high confidence, which in turn can assist LLMs in the process of in-context learning (Ram et al., 2023; Wang et al., 2023). To ensure the timeliness of knowledge, we utilize the daily updated search logs to repeat the above operation every day and construct daily separate indexes for the dual behavioral neighbors of queries and items. The neighbors could be retrieved from indexes for the corresponding query and item , denoted as and respectively. Need to clarify is that although we can obtain the behavior neighbors for the majority of queries or items, there is still a small portion of queries or items whose behavior neighbors cannot be obtained. In response to this situation, we will set the neighbors that cannot be obtained as empty in the prompt. Besides, we could obtain the attributes (i.e., brand, keyword, intent) of the query and item from the industrial knowledge base, which could be denoted as and . They can be treated as a supplement to the domain-specific information. Then we could construct our daily prompt based on the above information. The daily prompt will be send to LLMs for further operation.
4.2. Progressive Prompting and Aggregation
@@
4.2.1. Progressive Prompting
It is designed for making LLMs sensitive to the diverse aspects for the relevance judgement and improving the robustness of the model performance. Specifically, we firstly decompose the mentioned prompt above and construct least-to-most prompts step by step (Zhou et al., 2022; Wei et al., 2022) as:
@@ 4.2.1. Progressive Prompting
Then LLMs could reason the likelihood of the verbalizer exploiting the least-to-most prompts step by step for the sensitivity to relevance judgement and stable prediction results, which could be dubbed as:
Simply, all least-to-most prompts share the same relevance label and all least-to-most sub-tasks are supervised by the cross-entropy loss in parallel explicitly. The aforementioned method can be extended to progressive prompts and the supervised loss could be denoted as , where represents the loss of the sub-task corresponding to -th prompt.
4.2.2. Progressive Aggregation
Once we obtain the least-to-most probabilities, we expect to learn the progressive relationship of solutions to sub-tasks. Intuitively, Our least-to-most paradigm is incremental not only in terms of information volume, but also in importance. Hence, we adopt the kernel function to model the incremental tendency and the choice of kernel function can be a Gaussian kernel function, exponential kernel function, logarithmic decay kernel function, etc. After experimental comparison, we have chosen the exponential kernel function , where is a learnable parameter and is the degree of attenuation of the -th prompt defined by us. The overall relevance score could be acquired through aggregating the probabilities from the least-to-most sub-tasks progressively as:
@@ 4.2.2. Progressive Aggregation
5. Industrial Implementation for LLMs

### v2

@@
Abstract
††footnotetext: Correspondence to Jia Xu and Wei Zhang.Relevance modeling is a critical component for enhancing user experience in search engines, with the primary objective of identifying items that align with users’ queries. Traditional models only rely on the semantic congruence between queries and items to ascertain relevance. However, this approach represents merely one aspect of the relevance judgement, and is insufficient in isolation. Even powerful Large Language Models (LLMs) still cannot accurately judge the relevance of a query and an item from a semantic perspective. To augment LLMs-driven relevance modeling, this study proposes leveraging user interactions recorded in search logs to yield insights into users’ implicit search intentions. The challenge lies in the effective prompting of LLMs to capture dynamic search intentions, which poses several obstacles in real-world relevance scenarios, i.e., the absence of domain-specific knowledge, the inadequacy of an isolated prompt, and the prohibitive costs associated with deploying LLMs. In response, we propose ProRBP, a novel gressive etrieved ehavior-augmented rompting framework for integrating search scenario-oriented knowledge with LLMs effectively. Specifically, we perform the user-driven behavior neighbors retrieval from the daily search logs to obtain domain-specific knowledge in time, retrieving candidates that users consider to meet their expectations. Then, we guide LLMs for relevance modeling by employing advanced prompting techniques that progressively improve the outputs of the LLMs, followed by a progressive aggregation with comprehensive consideration of diverse aspects. For online serving, we have developed an industrial application framework tailored for the deployment of LLMs in relevance modeling. Experiments on real-world industry data and online A/B testing demonstrate our proposal achieves promising performance.
Towards Boosting LLMs-driven Relevance Modeling with Progressive Retrieved Behavior-augmented Prompting
Zeyuan Chen1, Haiyan Wu2, Kaixin Wu1, Wei Chen1, Mingjie Zhong1, Jia Xu1, Zhongyi Liu1, Wei Zhang 3 1Ant Group, 2Alibaba Group, 3School of Computer Science and Technology, East China Normal University {chenzeyuan, daniel.wkx, qianmu.cw, mingjie.zmj, steve.xuj, zhongyi.lzy}@antgroup.com, wuhaiyan.why@taobao.com, zhangwei.thu2011@gmail.com
1 Introduction
@@
To address the above problems, we propose a novel gressive etrieved ehavior-augmented rompting framework for integrating search scenario-oriented knowledge with LLMs (dubbed as ProRBP). To acquire domain-specific knowledge in time, we perform a user-driven behavior neighbor retrieval from the daily updated search logs, retrieving candidates that users consider to meet their expectations currently. Then we anticipate employing advanced prompting techniques that progressively improve the outputs of the LLMs, followed by a progressive aggregation with comprehensive consideration of diverse aspects to form a holistic relevance model. As for the online serving of LLMs, we design an industrial implementation framework enabling LLMs to fully handle search relevance scenarios with the affordable cost.
@@
We propose ProRBP with two novel plug-in modules. Firstly, a user-driven behavior neighbors retrieval is developed to acquire domain-specific knowledge in time. Secondly, the proposal of progressive prompting and aggregation can strengthen the judgement of relevance for LLMs.
We explore an industrial implementation A.1 enabling LLMs to fully handle search relevance scenarios in Alipay search with the affordable cost and latency.
We demonstrate ProRBP framework achieves superior performance through experiments on real-world industry data and online A/B testing. It has been deployed online and outperforms prior approaches (Chen et al., 2023) on core metrics.
2 Related Work
Relevance modeling in search can be viewed as a text matching problem as the sub-domain of information retrieval (IR). The majority of work focuses on distinctions at the semantic level, while a minority of methods judge the relevance from a behavioral perspective.
@@ 2 Related Work
In recent times, pre-trained models like BERT (Devlin et al., 2018) have made remarkable strides in Natural Language Understanding (NLU). As a result, representation-based (Yao et al., 2022; Lu et al., 2020; Reimers and Gurevych, 2019; Jin et al., 2023) and interaction-based architectures (Wang et al., 2019) are proposed to leverage the capabilities of these models to encode semantic correlations between queries and items. Most recently, Large Languages Models (LLMs) like GPT (Radford et al., 2019), LLaMA (Touvron et al., 2023), BLOOM (Scao et al., 2022) and GLM (Du et al., 2022) trained on massive corpora of texts have shown their superior ability in language understanding, generation, interaction, and reasoning tasks. (Sun et al., 2023) investigates the potential of utilizing LLMs for searching and demonstrates that appropriately instructs ChatGPT and GPT-4 can produce competitive and even superior results to supervised methods widely used information retrieval benchmarks. (Chen et al., 2023) tries to deal with long-tail query-item matching through LLMs efficiently and effectively. In these research work, they tend to exploit world knowledge stored in parameters of LLMs to judge the relevance between the query and item. However, general LLMs can not adapt to the industrial scenario due to the lack of domain-specific knowledge and insensitivity to the relevance judgement. In this work, we try to explore the LLMs-driven relevance modeling comprehensively in Alipay search engine to meet the above requirements.
@@ 3 Problem Formulation
4 Methodology
In this section, we elaborate the proposed ProRBP framework with its novel plug-in modules as depicted in Figure 2.
4.1 User-driven Behavior Neighbor Retrieval
@@
Generally, LLMs possess an exhaustive world knowledge with the benefit of massive corpora of texts pre-training. However, LLMs still struggle to understand user search intentions for short and ambiguous queries and items from Alipay search due to the lack of specialized and rapidly evolving domain knowledge. A promising approach for addressing the above is retrieval-augmented language modeling (Gao et al., 2023), grounding the LLMs during generation by conditioning on relevant candidates retrieved from an external knowledge source. Drawing inspiration from this, we devise a user-driven behavior neighbor retrieval module. This module could retrieve the daily search logs of users to obtain daily changing behavior neighbors that users consider relevant. The pipeline of user-driven behavior neighbor retrieval is depicted as Figure 1. Due to the limited number of items displayed for a query in Alipay search scenario, almost all results can be seen by users. With this in mind, we could analyze that a higher click-through rate reflects that users believe the corresponding query-item pairs can better meet their search intents and needs when the exposure PV (i.e., page view) reaches a certain quantity. Thus, we filter out the query-item pairs with less than 100 exposure PV and formulate the remaining search logs as high-confidence logs. Then, we utilize the high-confidence logs from the past month to calculate the click-through rate for the exposed query-item pairs. To mitigate the effect of noises, query-item pairs are selected above a click-through threshold (e.g., 0.2) and the neighbors are arranged in descending order based on click-through rate from the query and item perspective, respectively. And we only choose top-K (e.g., 20) neighbors for corresponding query and item. Through this approach, it is possible to select dual behavior neighbors with high confidence, which in turn can assist LLMs in the process of in-context learning (Ram et al., 2023; Wang et al., 2023; Liao et al., 2024; Liu et al., 2024; Wei et al., 2024). To ensure the timeliness of knowledge, we utilize the daily updated search logs to repeat the above operation every day and construct daily separate indexes for the dual behavioral neighbors of queries and items. The neighbors could be retrieved from indexes for the corresponding query and item , denoted as and respectively. Need to clarify is that though we can obtain the behavior neighbors for the majority of queries or items, there is still a small portion of queries or items whose behavior neighbors cannot be obtained. In response to this situation, we will set them as empty in the prompt. Besides, we could obtain the attributes (i.e., brand, keyword, intent) of the query and item from the industrial knowledge base, which could be denoted as and . They can be treated as a supplement to the domain-specific information. Then we could construct our daily prompt based on the above information. The daily prompt will be send to LLMs for further operation.
4.2 Progressive Prompting and Aggregation
@@
Progressive Prompting. It is designed for making LLMs sensitive to the diverse aspects for the relevance judgement and improving the robustness of the model performance. Specifically, we firstly decompose the mentioned prompt above and construct least-to-most prompts step by step (Zhou et al., 2022; Wei et al., 2022) as:
@@
Then LLMs could reason the likelihood of the verbalizer exploiting the least-to-most prompts step by step for the sensitivity to relevance judgement and stable prediction results. Simply, all least-to-most prompts share the same relevance label and all least-to-most sub-tasks are supervised by the cross-entropy loss in parallel explicitly. The aforementioned method can be extended to progressive prompts and the supervised loss could be denoted as , where represents the loss of the sub-task corresponding to -th prompt.
Progressive Aggregation. Once we obtain the least-to-most probabilities, we expect to learn the progressive relationship of solutions to sub-tasks. Intuitively, Our least-to-most paradigm is incremental not only in terms of information volume, but also in importance. Hence, we adopt the kernel function to model the incremental tendency and the choice of kernel function can be a Gaussian kernel function, exponential kernel function, logarithmic decay kernel function, etc. After experimental comparison, we have chosen the exponential kernel function , where is a learnable parameter and is the degree of attenuation of the -th prompt defined by us. The overall relevance score could be acquired through aggregating the probabilities from the least-to-most sub-tasks progressively as:
@@
5 Experiments
5.1 Experimental Setup
Datasets. In order to evaluate the performance of all the models with reliability, we select the real-world industry data used in mini apps search scenario of Alipay search engine and present its statistics in Table 1. The dataset is labeled by human annotators, where Good and Bad annotations denote label 1 and 0 respectively. User historical behavior data is sampled from the search logs of the search engine. Although datasets such as WANDS***https://github.com/wayfair/WANDS/tree/main and MSLR†††https://www.microsoft.com/en-us/research/project/mslr/ are publicly available, they do not contain the requisite user historical behavior data. Hence, we select this in-house data to evaluate the proposed framework. Other related work (Yao et al., 2022, 2021; Zhu et al., 2021; Li et al., 2021; Chen et al., 2023) also selects one in-house data to evaluate the proposed approaches. The partial data of this paper was released before‡‡‡https://github.com/alipay/BehaviorAugmentedRelevanceModel.
Baseline Models. We select a set of popular NLU (Natural Language Understanding)-based, behavior-based and NLG (Natural Language Generation)-based relevance models as baselines. For the NLU-based models, we choose three common models including two-tower and single-tower architectures: DSSM (Shen et al., 2014), ReprBert (Yao et al., 2022), Bert (Devlin et al., 2018). The behavior-based models all consider constructing the pre-training dataset (MASM (Yao et al., 2021)) or the topology structure based on click behaviors (TextGNN (Zhu et al., 2021), AdsGNN (Li et al., 2021), BARL-ASe (Chen et al., 2023)). For the NLG-based models, we choose two foundation models including two set of architectures: the causal decoder (BLOOM (Scao et al., 2022)) and the prefix decoder (GLM (Du et al., 2022)).
| Dataset | # Sample | # Query | # Item | # Good | # Bad |
| Train | 773,744 | 87,499 | 88,724 | 460,610 | 313,134 |
| Valid | 97,032 | 40,754 | 25,192 | 57,914 | 39,118 |
| Test | 96,437 | 40,618 | 25,004 | 57,323 | 39,114 |
Evaluation Metrics. We use Area Under Curve (AUC), F1-score (F1), and False Negative Rate (FNR) to measure the multidimensional performance of all models. AUC and F1 are commonly used in the studied area, of which higher metric values represent better model performance. Conversely, lower False Negative Rate (FNR) values are preferable, as they indicate a lower false filtering rate of models. Note that AUC often serves as the most significant metric in our task while the others provide auxiliary supports for our analysis.
Model Implementations. We tune our model for 5 epochs with batch size 64 and learning rate 3e-05 in the supervised fine-tuning stage of LLMs. For the foundation models, we select the 1.1B BLOOM§§§https://huggingface.co/bigscience/bloom-1b1 and GLM with different magnitude of parameters (e.g., 0.3B, 2B and 10B) pre-trained by Alipay. The number of retrieved behavior neighbors is set to 20. We tune the parameter within the ranges of {0, 0.05, 0.1, 0.2, 0.3, 0.5, 1.0}. We conduct the experiments on NVIDIA Tesla A100 GPUs.
| Method | AUC | F1 | FNR (-) |
| DSSM | 0.8356 | 0.8210 | 0.1389 |
| ReprBert | 0.8388 | 0.8376 | 0.1280 |
| Bert | 0.8540 | 0.8534 | 0.1150 |
| MASM | 0.8547 | 0.8318 | 0.1289 |
| TextGNN | 0.8847 | 0.8489 | 0.1290 |
| AdsGNN | 0.8878 | 0.8458 | 0.1454 |
| BARL-ASe | 0.9078 | 0.8658 | 0.1054 |
| GLM-0.3B | 0.8608 | 0.8585 | 0.1106 |
| +ProRBP | 0.9105 | 0.8778 | 0.1035 |
| BLOOM-1.1B | 0.8543 | 0.8511 | 0.1174 |
| +ProRBP | 0.8991 | 0.8675 | 0.1121 |
| GLM-2B | 0.8619 | 0.8598 | 0.1065 |
| +ProRBP | 0.9120 | 0.8776 | 0.1007 |
| GLM-10B | 0.8751 | 0.8656 | 0.1006 |
| +ProRBP | 0.9143 | 0.8801 | 0.0973 |
5.2 Experimental Results
Performance Comparison. Table 2 shows the overall comparison with baselines. Our findings indicate that DSSM performs poorly since it merely encodes the embeddings of the query and item independently. By further comparing DSSM with ReprBert, we find the performance is improved to a certain degree. This demonstrates the pre-trained models utilizing the large corpus can bring additional gains. Compared to ReprBert, Bert achieves better performance, as the interaction-based models possess more advanced text relevance modeling abilities than representation-based models.
For the models of MASM, TextGNN, and AdsGNN, they exploit historical behavior data to build pre-training dataset or click graph to enhance the effectiveness of search systems. And they achieve significantly better results than the above-mentioned methods in the metric of AUC rather than F1 and FNR. This can be attributed to the introduction of auxiliary signals, which may bring improvement for the ranking ability of models but introduce some noises leading to the loss of F1 and FNR. The performance differences among them depend on the utilization of behavior data and modeling granularity. The newly proposed model BARL-ASe achieves the best performance in behavior-based relevance models.
GLM-0.3B, BLOOM-1.1B, GLM-2B and GLM-10B denote that we use the corresponding language models with the amount of parameters to perform relevance judgements exploiting the naive prompts as stated in Section 3. They exhibit relatively good performance, demonstrating the positive effect of massive training datasets and large-scale parameters. Performance variances among them with different parameters appear minimal, possibly because the simplicity of the relevance task doesn’t fully tap into the advantages of LLMs. Their performance has a certain gap compared to behavior-based models especially on the core metric AUC. This may be attributed to the lack of domain-specific knowledge and the limitations of an isolated prompt. Thus, ProRBP framework is proposed to address the mentioned problem and significantly improve the foundation models’ performance. The performance differences between BLOOM and GLM may stem from differences in architecture and training corpora.
Overall, exploiting our framework ProRBP could yield the best performance and ensures efficiency and economy. In the online serving A.1, we select GLM-10B to perform offline inference and GLM-2B to proceed online prediction considering affordable cost and efficiency. Besides, they both obtain better gains under all the metrics on the evaluation dataset compared to the second-best performed model BARL-ASe that was deployed in our scenario before.
| Method | AUC | F1 | FNR (-) |
| ProRBP | 0.9120 | 0.8776 | 0.1007 |
| -BNR | 0.8943 | 0.8656 | 0.1037 |
| -PPA | 0.8872 | 0.8645 | 0.1045 |
| -Both | 0.8619 | 0.8598 | 0.1065 |
Ablation Study. To investigate the contributions of key components, we provide the following variants of our complete framework: (1) “-BNR” denotes discarding user-driven behavior neighbor retrieval strategy; (2) similarly, “-PPA” denotes discarding progressive prompting and aggregation module. (3) “-Both” represents removing two modules mentioned above.
The results shown in Table 3 lead to the conclusion that both “BNR” and “PPA” strategies yield significantly positive results. And “PPA” demonstrates greater significance than “BNR”.
5.3 Parameter Sensitivity
From the variation trends of Figure 3, we could observe that: (1) As the number of behavioral neighbors increases, the model’s performance continues to improve, tending to plateau at around 20. (2) The exponential kernel function yields the best result among the mean-pooling (Mean), Gaussian (Norm) and logarithmic (Log) kernel functions. (3) The better results are achieved when setting to a suitable value. Too large (e.g., 1.0) or small (e.g., 0.01) value will lead to a performance drop.
5.4 Online A/B Testing
The proposed method has been deployed in the relevance stage of Alipay search platform providing search service of mini apps and demonstrated its significant performance gains in online A/B testing compared with the previous model BARL-ASe. The each experiment takes about 3% proportion of Alipay search traffic for two weeks. Compared with the previously deployed model, the proposed method improves the valid PV-CTR¶¶¶the number of valid clicks divided by the number of searches by 0.33% on average without causing an increase in latency. And the results of human annotations show the model can reduce the rate of irrelevant results by 1.07% points on average. The results demonstrate that our proposal can improve the experience of users.
6 Conclusion
This paper studies the relevance modeling problem by integrating world knowledge stored in the parameters of LLMs with specialized domain knowledge represented by user behavior data for achieving promising performance. The novel framework ProRBP is proposed, which innovatively develops user-driven behavior neighbor retrieval module to learn domain-specific knowledge in time and introduces progressive prompting and aggregation module for considering diverse aspects of the relevance and prediction stability. We explore an industrial implementation to deploy LLMs to handle full-scale search traffics of Alipay with acceptable cost and latency. The comprehensive experiments on real-world industry data and online A/B testing validate the superiority of our proposal and the effectiveness of its main modules.
Appendix A Appendix
A.1 Industrial Implementation for LLMs